# Hypergraph adjusted plus-minus

### v2

@@ Abstract
In team sports, traditional ranking statistics do not allow for the simultaneous evaluation of both individuals and combinations of players. Metrics for individual player rankings often fail to consider the interaction effects between groups of players, while methods for assessing full lineups cannot be used to identify the value of lower-order combinations of players (pairs, trios, etc.). Given that player and lineup rankings are inherently dependent on each other, these limitations may affect the accuracy of performance evaluations. To address this, we propose a novel adjusted box score plus-minus (APM) approach that allows for the simultaneous ranking of individual players, lower-order combinations of players, and full lineups. The method adjusts for the complete dependency structure and is motivated by the connection between APM and the hypergraph representation of a team. We discuss the similarities of our approach to other advanced metrics, demonstrate it using NBA data from 2012-2022, and suggest potential directions for future work.
@@ 1 Introduction
One of the most common approaches to measuring an individual’s effect in team sports is box-score plus-minus (PM). Originally used in hockey, a player’s PM is their team’s cumulative difference in score when that player is in the game. While intuitively appealing, PM is sensitive to substitution patterns. For instance, one very good player can artificially inflate the PM of their teammates, and quantifying this effect becomes challenging when certain combinations of players consistently share playing time. The most common approach to account for this is known as adjusted PM (APM). APM was developed for basketball (Rosenbaum, 2004) and hockey (Macdonald, 2011) as a method for comparing individual players while accounting for those that play together. We provide the mathematical details of APM in Section 3.
The intention of APM and its variants (Hvattum, 2019) is to isolate an individual player’s effect. However, this approach fails to capture any interaction effects between players. Experience, complementary skill sets, and general synergy often result in a group of players whose combined effect surpasses the sum of their individual contributions. This phenomena is in part why substitution patterns are not uniformly random, and why managers and coaches play a crucial role in the success of a team. Likewise, when assessing and ranking an individual’s contribution to a team’s success, it is critical to account for the player’s ability to create synergies with the various combinations of their teammates. We refer to these various combinations (individuals, pairs, trios, etc.) as generalized lineups and propose a novel approach to quantify the contribution of all generalized lineups for a team simultaneously.
@@
Both of these approaches have obvious drawbacks, and neither is able to rank the generalized lineups while accounting for interaction effects. To fill the gap between these two extreme solutions, we introduce a new class of methods for ranking lineups and players that adjusts not only for a full lineup, but all lower-order combinations of players. Our methods are motivated by the hypergraph representation of a lineup, which we show is a generalization of the design matrix used in the ordinary APM method. The result is a unified approach for simultaneously ranking any generalized lineup including individuals, pairs, and full lineups. Furthermore, our methods lend themselves to interpretable visualizations that allow one to quickly identify positive and negative synergies among groups of players on a team.
The rest of the paper is organized as follows. In Section 2, we review related methods in the literature for ranking lineups. We then provide the mathematical background for APM, used to rank individual players, in Section 3. In doing so, we propose a new variant, pairs APM, which is able to rank and adjust for pairs of players, but is ultimately too computationally expensive for higher-order generalized lineups. In Section 4, we review the necessary graph theory terminology and notation, and then make the connection between APM and hypergraphs. This leads to the development of our two new ranking methods: hypergraph APM and line graph APM. We apply these new methods to 10 years of NBA data from 2012-2022 in Section 5 and conclude with a discussion in Section 6 of our methods, their limitations, and potential future directions.
@@ 2 Related Work
The method most similar to ours is Devlin and Uminsky (2020), which takes a spectral approach. The authors define a success function $f(L)=\text{pm}_{L}$, where $L$ is one of the $\binom{15}{5}$ 5-person lineups from a roster of 15 and $\text{pm}_{L}$ is the lineup’s plus-minus. The success function is shown to decompose orthogonally into the generalized lineups, hence providing a natural measure of contribution from each generalized lineup. Inference is performed by relating the success function to the Johnson graph. To account for the number of possessions a lineup plays together, the authors normalize their results by the log of the number of possessions played by each lineup, which they call spectral contribution per log possession (SCLP). There is not a natural means to adjusting for opponent with SCLP.
@@ 3 Adjusted plus-minus
To illustrate APM, imagine a team with 5 players labeled A-E, and a 10-minute game with lineups of size 3. The results of one such game are summarized in Table 1. In the linear model described above, the player columns (A:E) define the design matrix $X$ and ‘Score Dif’ is the response vector $Y$. We see that the first lineup (A, B, C) was the most successful with a score differential of +3, indicating that this lineup outscored the other team by 3 total points in the 3 minutes that A, B, and C played together. We also see that Player A has the top ranking PM score of +5, which can be computed via the dot product $X_{\cdot,A}\cdot Y$. However, Player A always played with Player B, so it is unclear which of the two players deserves more credit for their lineups’ successes.
@@
In our example, the fitted coefficients from regressing $Y$ on $X$, shown in Table 2, result in the following player rankings (in order from best to worst): A, C, B = D, E. This model fits the data perfectly given $n=p=5$, and the coefficient estimates are very sensitive to small perturbations in the data. Furthermore, the existence of rank ties, which we see with players D and E, is undesirable.
@@
More advanced APM metrics incorporate both home and away players, use a penalized regression setup to stabilize the estimates (Sill, 2010), and incorporate playing time and additional box score statistics (Hvattum, 2019). Continuing with our prior example, we can employ a penalized regression scheme that prevents overfitting, and include a diagonal weight matrix of Time, $W$, that takes into account the total time each lineup was active. For instance, ridge regression (Hoerl and Kennard, 1970) yields the APM solutions $\widehat{\beta_{r}}=(X^{T}WX+\lambda I)^{-1}X^{T}WY$. In our example, after using cross-validation to choose an appropriate value of the penalization parameter $\lambda$, this results in the improved ranking scheme: A, C, B, D, E. Note that each $\widehat{\beta}$ is an estimated parameter, so we could quantify the uncertainty and provide the implied rank intervals. We omit this here, for clarity, but discuss it later in Section 5 with our application to real data.
While a popular player-ranking metric, a shortcoming of APM is that it does not adjust for generalized lineups. Consider the scenario where two players excel when paired together but perform less effectively when playing separately. APM fails to capture this nuance, and may result in both individual players being ranked close to the median depending on their frequency of playing together versus apart. On the other hand, some individually great players may not play well together if, for example, their skill sets overlap. It is therefore necessary to account for combinations of players within full lineups.
@@ 3.1 Pairs APM
We include the coefficients from PAPM in Table 2 found using weighted ridge regression. Interestingly, the individual player rankings changed, compared to the $\hat{\beta_{r}}$ estimates, e.g. player B now outranks player C. Furthermore, there is the added benefit of ranking pairs by utilizing the regression coefficients associated with the additional columns in the design matrix. For context, the pair rankings from best to worst are AB, BC, AC, AD, BD, and so on.
@@ 4.1 Background
Graphs, or networks, are relational data objects composed of nodes or vertices that are linked by edges. Edges represent various types of relationships or interactions between nodes, and a variety of statistical methods can be utilized to model and understand the complex dependencies captured by the network (Kolaczyk and Csárdi, 2014). Given that team lineups are inherently interconnected, networks of game/team data are a natural representation to consider when ranking generalized lineups.
Mathematically, a network is a graph $G=(V,E)$ with node set $V=\{1,2,3,...,n\}$ and edge set $E~{}\subset~{}V\times V.$ Networks are often represented by an adjacency matrix $A\in\mathbb{R}^{n\times n}$ where $A_{ij}~{}=~{}1$ whenever $(i,j)\in E$ and 00 otherwise. Another important matrix representation is the graph Laplacian $L=D-A$ where $D$ is the diagonal degree matrix, i.e. $D_{ii}=\sum_{j\in V}A_{ij}$. Finally, the incidence matrix $M$ has a row for each node, a column for each edge in the network, and $M_{ij}=1$ if node $i$ is incident to edge $e_{j}$.
Importantly, we can realize the design matrix described above in Section 3 as the transpose of the incidence matrix of a hypergraph. A hypergraph is a generalization of a graph where edges can connect more than two vertices. In this interpretation, the nodes of the hypergraph are the players and edges represent lineups. This connection between APM and hypergraphs motivates a deeper exploration into the possibility of a network analysis to inform the ranking of players and generalized lineups.
To date, statistical methods leveraging network information via hypergraphs are less developed than those for simple graphs. This suggests a transformation of the hypergraph would be helpful. One such approach is to transform the edge-node relationship of the hypergraph via its line graph. A line graph is itself a graph, and is derived from another graph by associating vertices of the line graph with edges of the original graph. For the line graph of a hypergraph, each hypergraph edge is a node, and two nodes in the line graph are connected if their corresponding hypergraph edges share at least one node.
The edges of the line graph can be weighted based on the number of common nodes between the hypergraph edges. Henceforth, we utilize the Jaccard similarity,
@@
where $e_{i}$ is the set of nodes incident to hypergraph edge $i$. The hypergraph induced by the lineup data from Table 1 is illustrated in Figure 1, along with its line graph. We see, for example, that node ABD in the line graph – representing the hypergraph edge and lineup ABD – is connected to node ABE by an edge with weight $\frac{2}{4}$, since out of the combined set of 4 players A, B, D and E, only players A and B are common to both nodes. The score differentials $Y$ are associated with a node in the line graph, and the edges represent a level of similarity between lineups. All of the notation introduced in the previous paragraph follows, however the binary adjacency matrix is replaced by its weighted counterpart, such that $A_{ij}=w_{ij}$ whenever $(i,j)\in E$.
Figure 1: An example hypergraph (left) and its corresponding line graph (right). The incidence matrix of the hypergraph is exactly the transpose of the design matrix in the lineup data in Table 1.
The hypergraph and its corresponding line graph, as currently described, are based solely on full lineups. This is a $k$-regular hypergraph, where each hyperedge connects exactly $k=3$ nodes. However, we can expand our representation to include lower-order player combinations by introducing additional hypergraph edges and their associated nodes in the line graph. We refer to this as the extended hypergraph.
Continuing with our toy example, we add additional rows to Table 1 describing the aggregate pair and solo player results, as shown, in the case of Player A, in Table 4. Figure 2 pictures the extended hypergraph and its line graph. Note that the line graph now has a node for all generalized lineups, and generalized lineups are connected via a weighted edge based on Jaccard similarity.
@@
Table 4: Extended lineup data from Table 1 with generalized lineups. In this case, we show the additional rows corresponding to all lineups containing player A.
Figure 2: The extended hypergraph defined by Table 4 and its corresponding line graph.
@@ 4.2.1 Hypergraph APM
As an alternative to PAPM that is scalable for all generalized lineups, we introduce hypergraph adjusted plus-minus (HAPM). HAPM is the same multiple linear regression described in Section 3 for APM, but with the design matrix replaced by the transpose of the incidence matrix of the extended hypergraph. In this case, the columns of the design matrix represent each player, the rows are combinations of players, and the entry is 1 if that player is in that row’s combination of players. That is, $X_{\text{HAPM}}$ is the new design matrix with $({X_{\text{HAPM}}})_{ij}=1$ if player $j$ is in the combination of players represented in row $i$ and 0 otherwise. Note that if we restrict the rows to full lineups only, then this is exactly ordinary APM and the design matrix is the transpose of the incidence matrix of the $k$-regular (non-extended) hypergraph. As before, the response is the plus-minus for a given row and, using ridge regression, we obtain coefficients $\widehat{\beta}_{\text{HAPM}}$ at the individual player level that consider all generalized lineups included in the design matrix.
@@ 4.2.2 Line graph APM
HAPM utilizes the hypergraph representation implicitly, yet it does not explicitly consider the connectivity patterns within the hypergraph. To address this, we draw inspiration from functional data analysis (Ramsay and Silverman, 2002) and propose a Bayesian regression approach. This approach resembles kernel regression on a graph (Smola and Kondor, 2003), allowing us to extend regression-based techniques to graph-indexed data, where conventional models and assumptions may not be applicable. In our case, it allows for a ranking of individual players and generalized lineups while considering all active combinations of players on a team.
Utilizing the extended line graph representation of a team, we view the plus-minus associated with each node as a random response variable $Y_{v}$. Our goal is to predict this value using information encoded in the network, and then rank the players and the generalized lineups based on these predictions. We use a Bayesian regression model that is specified as follows:
@@
Our choice of the graph Laplacian is designed to exploit information in the network to inform the model and force the predicted values to be smooth with respect to the topology of the network. That is, nodes are assigned values based on their connectivity in the network and adjacent nodes will have similar predicted values. In fact, the log prior can be expanded as
@@
showing that the term penalizes the weighted sum of squared differences for coefficients of adjacent vertices in the network (Kolaczyk and Csárdi, 2014). Moreover, pairs of nodes with larger edge weights, or more similar generalized lineups, will be forced to have more similar predicted values. Given that $L_{w}$ is a symmetric matrix, it realizes the eigendecomposition $L_{w}=\Phi\Xi\Phi$, and we can therefore represent $\beta$ using a basis expansion with respect to the eigenvectors of $L_{w}$ by letting $\beta=\Phi\theta$ (Upton and Carvalho, 2017). In short, we use the eigenvectors of the Laplacian as predictor variables, find a set of corresponding coefficients $\hat{\theta}$ through Bayesian linear regression, and then perform matrix-vector multiplication to solve for the vertex-indexed intercepts $\widehat{\beta_{v}}=\widehat{Y_{v}}.$
We do not have to use the full set of eigenvectors to represent $\beta$. Instead, we can extract most of the information in $L_{w}$ through a low-order representation such that $\beta=\Phi_{1:\tau}\theta$, where $\Phi_{1:\tau}$ represents the first $\tau$ eigenvectors (ordered by the eigenvalues $\xi_{1}<\dots<\xi_{\tau}$). In practice, we select $\tau$ based on an ‘elbow’ method, examining sequential differences in the size of ${\xi_{i}}$.

### v3

@@ Abstract
In team sports, traditional ranking statistics do not allow for the simultaneous evaluation of both individuals and combinations of players. Metrics for individual player rankings often fail to include the interaction effects between groups of players, while methods for assessing full lineups cannot be used to identify the value of lower-order combinations of players (pairs, trios, etc.). Given that player and lineup rankings are inherently dependent on each other, these limitations may affect the accuracy of performance evaluations. To address this, we propose a novel adjusted plus-minus (APM) approach that allows for the simultaneous ranking of individual players, lower-order combinations of players, and full lineups within a team. The method adjusts for the complete dependency structure and is motivated by the connection between APM and the hypergraph representation of a team. We discuss the similarities of our approach to other advanced metrics, demonstrate it using NBA data from 2012-2022, and suggest potential directions for future work.
@@ 1 Introduction
One of the most common approaches to measuring an individual’s effect in team sports is raw plus-minus (PM). Originally used in hockey, a player’s PM is their team’s cumulative difference in score when that player is in the game. While intuitively appealing, PM is sensitive to substitution patterns. For instance, one very good player can artificially inflate the PM of their teammates, and quantifying this effect becomes challenging when certain combinations of players consistently share playing time. The most common approach to account for this is known as adjusted PM (APM). APM was developed for basketball (Rosenbaum, 2004) and hockey (Macdonald, 2011) as a method for comparing individual players while accounting for those that play together. It has also been extended, via a Bayesian framework, to soccer and American football (Matano et al., 2023; Sabin, 2021). We provide some mathematical details of APM in Section 3.
The intention of APM and its variants is to isolate an individual player’s effect (Hvattum, 2019). However, this approach fails to capture any interaction effects between players. Experience, complementary skill sets, and general synergy often result in a group of players whose combined effect surpasses the sum of their individual contributions. This phenomena is in part why substitution patterns are not uniformly random, and why managers and coaches play a crucial role in the success of a team. Likewise, when assessing and ranking an individual’s contribution to a team’s success, it is critical to account for the player’s ability to create synergies with the various combinations of their teammates. We refer to these various combinations (individuals, pairs, trios, etc.) as generalized lineups and propose a novel approach to quantify the contribution of all generalized lineups for a team simultaneously.
@@
Both of these approaches have obvious drawbacks, and neither is able to rank the generalized lineups while accounting for interaction effects. To fill the gap between these two extremes, we propose a solution for ranking lineups and players that adjusts not only for a full lineup, but all lower-order combinations of players. Our methods are motivated by the hypergraph representation of a lineup, which we show is a generalization of the design matrix used in the ordinary APM method. The result is a unified approach for simultaneously ranking any generalized lineup including individuals, pairs, and full lineups. Furthermore, our methods lend themselves to interpretable visualizations that allow one to quickly identify positive and negative synergies among groups of players on a team.
The rest of the paper is organized as follows. In Section 2, we review related methods in the literature for ranking lineups. We then provide the mathematical background for APM, used to rank individual players, in Section 3. In doing so, we discuss pairs APM (a variant of APM that can be used to rank and adjust for pairs of players, but is ultimately too computationally expensive for higher-order generalized lineups). In Section 4, we review the necessary graph theory terminology and notation, and then make the connection between APM and hypergraphs. This leads to the development of our two new ranking methods. We apply these new methods to 10 years of NBA data from 2012-2022 in Section 5 and conclude with a discussion in Section 6 of our methods, their limitations, and potential future directions.
@@ 2 Related Work
The method most similar to ours is Devlin and Uminsky (2020), which takes a spectral approach. The authors define a success function $f(L)=\text{pm}_{L}$, where $L$ is one of the $\binom{15}{5}$ 5-person lineups from a roster of 15 and $\text{pm}_{L}$ is the lineup’s raw plus-minus. The success function is shown to decompose orthogonally into the generalized lineups, hence providing a natural measure of contribution from each generalized lineup. Inference is performed by relating the success function to the Johnson graph. To account for the number of possessions a lineup plays together, the authors normalize their results by the log of the number of possessions played by each lineup, which they call spectral contribution per log possession (SCLP). There is not a natural means to adjusting for opponent with SCLP.
@@ 3 Adjusted Plus-Minus
To illustrate APM, we introduce a hypothetical example that will be referred to throughout the manuscript. Imagine a team with 5 players labeled A-E, and a 10-minute game with lineups of size 3. The results of one such game are summarized in Table 1. In the linear model described above, the player columns (A:E) define the design matrix $X$ and ‘Score Dif’ is the response vector $Y$. We see that the first lineup (A, B, C) was the most successful with a score differential of +3, indicating that this lineup outscored the other team by 3 total points in the 3 minutes that A, B, and C played together. We also see that Player A has the top ranking PM score of +5, which can be computed via the dot product $X_{\cdot,A}\cdot Y$. However, Player A always played with Player B, so it is unclear which of the two players deserves more credit for their lineups’ successes.
@@
In our example, the fitted coefficients from regressing $Y$ on $X$, shown in Table 2, result in the following player rankings (in order from best to worst): A, C, B = D, E. This model fits the data perfectly given $n=p=5$, and the coefficient estimates are very sensitive to small perturbations in the data. Furthermore, the existence of rank ties, which we see with players D and E, is undesirable. More advanced APM metrics incorporate both home and away players (i.e. each unique observation is a lineup of one team vs the other team), use a penalized regression setup to stabilize the estimates (Sill, 2010), and incorporate playing time and additional box-score statistics or external ratings (Hvattum, 2019). In our example, after using cross-validation to choose an appropriate level of regularization, a weighted ridge regression results in the improved ranking scheme: A, C, B, D, E (see $\widehat{\beta_{r}}$ in Table 2). Alternatively, if information was known a priori about our example players, we could recast the ridge regression into a Bayesian framework and incorporate this additional player information in the prior distribution (Matano et al., 2023).
@@
While a popular player-ranking metric, a shortcoming of APM and its variants is that it does not adjust for generalized lineups. Consider the scenario where two players excel when paired together but perform less effectively when playing separately. APM fails to capture this nuance, and may result in both individual players being ranked close to the median depending on their frequency of playing together versus apart. On the other hand, some individually great players may not play well together if, for example, their skill sets overlap. It is therefore necessary to account for combinations of players within full lineups.
@@ 3.1 Pairs APM
We include the coefficients from PAPM in Table 2 found using weighted ridge regression. With a new estimation procedure, the individual player rankings can change compared to the $\hat{\beta_{r}}$ estimates, e.g. player B now outranks player C (although this case may be due to the fact that, as a consequence of ridge regression, player B receives a larger contribution simply by virtue of appearing in four of the five lineups). Furthermore, there is the added benefit of ranking pairs by utilizing the regression coefficients associated with the additional columns in the design matrix. For context, the pair rankings from best to worst are AB, BC, AC, AD, BD, and so on.
@@ 4.1 Background
Graphs, or networks, are relational data objects composed of nodes (also known as vertices) that are linked by edges. Edges represent various types of relationships or interactions between nodes, and a variety of statistical methods can be utilized to model and understand the complex dependencies captured by the network (Kolaczyk and Csárdi, 2014). Given that team lineups are inherently interconnected, networks of game/team data are a natural representation to consider when ranking generalized lineups.
Mathematically, a network is a graph $G=(V,E)$ with node set $V=\{1,2,3,...,n\}$ and edge set $E~{}\subset~{}V\times V.$ Networks are often represented by an adjacency matrix $A\in\{0,1\}^{n\times n}$ where $A_{ij}~{}=~{}1$ whenever $(i,j)\in E$ and 00 otherwise. Other important matrix representations include:
  • the graph Laplacian: $L=D-A$, where $D$ is the diagonal degree matrix, i.e. $D_{ii}=\sum_{j\in V}A_{ij}$
  • the incidence matrix: $M,$ which has a row for each node, a column for each edge in the network, and $M_{ij}=1$ if node $i$ is incident to edge $e_{j}.$
A hypergraph extends the concept of a graph by allowing edges, called hyperedges, between more than two nodes. Hypergraphs are a natural representation for team lineups, where the nodes are players, and hyperedges represent lineups. The hypergraph visualization for our toy example is in the center of Figure 1. The nodes represent the players A-E, and are connected to each other by hyperedges that represent the lineups. We have colored the hyperedges for a clearer visualization, e.g. red (R) represents the lineup A-B-C, which appears in the first column of Table 1.
Importantly, we realize the design matrix in Table 1 as the transpose of the incidence matrix of its hypergraph. Figure 1 illustrates this connection between the design matrix and the hypergraph. This connection between APM and hypergraphs motivates a deeper exploration into the possibility of a network analysis to inform the ranking of players and generalized lineups.
[…]
Figure 1: The incidence matrix $M$ (left) for the hypergraph (middle) is exactly the transpose of the design matrix $X$ for APM given in Table 1. The hyperedges are colored to enhance the visualization: R (red), Y (yellow), G (green), B (blue), and P (purple). The line graph (right) of the hypergraph (middle) is formed by swapping the roles of nodes and hyperedges.
Statistical methods that incorporate network information through hypergraphs are less developed than those for standard graphs. A common approach to integrating the hypergraph structure into analysis is to transform the hypergraph into a standard graph. One such transformation is the line graph of the hypergraph.
A line graph is constructed by treating each hyperedge in the original hypergraph as a node in the new graph. Two nodes in the line graph are connected by an edge if their corresponding hyperedges in the hypergraph share at least one node. This transformation allows us to analyze relationships between hyperedges in a more conventional graph-based framework. To quantify the strength of these connections, we assign weights to the edges in the line graph. Specifically, we use Jaccard similarity to measure the overlap between two hyperedges. The Jaccard similarity between two hyperedges $e_{i}$ and $e_{j}$ is given by:
@@
where $e_{i}$ and $e_{j}$ are the set of nodes the belong to the corresponding hyperedges. For example, in our toy dataset (illustrated in Figure 1), consider the hyperedges ABD (yellow) and ABE (purple) in the hypergraph. In the corresponding line graph, these hyperedges become nodes, which are connected by an edge. The weight of this edge is calculated using Jaccard similarity: out of the combined set of 4 players A, B, D and E, players A and B are common to both nodes, resulting in an edge weight of $\frac{2}{4}.$ This weighting scheme reflects the degree of similarity between different hyperedges. Since we are working with a weighted graph, we replace the binary adjacency matrix with its weighted counterpart, such that $A_{ij}=w_{ij}$ whenever $(i,j)\in E$. Furthermore, note that by utilizing the line graph representation, the score differentials, $Y,$ are now associated with the nodes in the line graph.
The hypergraph and its corresponding line graph, as currently described, are based solely on full lineups. This is known as a $k$-regular hypergraph, where each hyperedge connects exactly $k=3$ nodes. However, we can expand our representation to include lower-order player combinations by introducing additional hyperedges. We refer to this as the extended hypergraph, and the transpose of its incidence matrix as the extended APM design matrix. Continuing with our toy example, we add additional rows to Table 1 describing the aggregate pair and solo player results, as shown in the case of Player A, in Table 4. Note that we do not include redundant rows in the matrix, i.e. the A-B pairing and score differential is not duplicated in the extended lineup data for the B-A pairing. We also create the line graph corresponding to the extended hypergraph; both are shown in Figure 2. In Figure 2, the line graph (right) has a node for each generalized lineup, and the generalized lineups are connected via a weighted edge based on Jaccard similarity.
@@
Table 4: Extended lineup data from Table 1 with generalized lineups. Here, we only show the additional rows corresponding to all lineups containing player A.
Figure 2: The extended hypergraph (left) defined by Table 4 and its corresponding line graph (right).
@@ 4.2 Methods
Here, we present our two new methods that rank generalized lineups that are adjusted for the full lineup and all lower-order combinations of players: the Hypergraph Adjusted Plus-Minus (HAPM) and the Line Graph Adjusted Plus-Minus (LAPM). In short, HAPM is regression using the extended APM design matrix and LAPM employs Bayesian regression using the line graph of the extended hypergraph with a prior that enforces smoothness on the fitted values of similar lineups.
@@ 4.2.1 Hypergraph APM
As an alternative to PAPM that is scalable for all generalized lineups, we introduce HAPM. HAPM is the same weighted ridge regression described in Section 3 for APM, but with the original design matrix replaced by the transpose of the incidence matrix of the extended hypergraph. In this case, the columns of the design matrix represent each player, the rows are combinations of players, and the entry is 1 if that player is in that row’s combination of players. That is, $X_{\text{HAPM}}$ is the new design matrix with $({X_{\text{HAPM}}})_{ij}=1$ if player $j$ is in the combination of players represented in row $i$ and 0 otherwise. Note that if we restrict the rows to full lineups only, then this is exactly ordinary APM (calculated at the team level) and the design matrix is the transpose of the incidence matrix of the $k$-regular (non-extended) hypergraph. As before, the response is the raw plus-minus for a given row. Using ridge regression, we obtain coefficients $\widehat{\beta}_{\text{HAPM}}$ at the individual player level that account for all generalized lineups included in the design matrix.
@@ 4.2.2 Line Graph APM
HAPM utilizes the hypergraph representation implicitly, yet it does not explicitly consider the connectivity patterns within the hypergraph. To address this, we draw inspiration from functional data analysis (Ramsay and Silverman, 2002) and propose a Bayesian regression approach using the line graph of the extended hypergraph. This approach resembles kernel regression on a graph (Smola and Kondor, 2003), allowing us to extend regression-based techniques to graph-indexed data, where conventional models and assumptions may not be applicable. In our case, it allows for a ranking of individual players and generalized lineups while considering all active combinations of players on a team.
Utilizing the extended line graph representation of a team, we view the raw plus-minus associated with each node as a random response variable $Y_{v}$. Our goal is to predict this value using information encoded in the network, and then rank the players and the generalized lineups based on these predictions. We use a Bayesian regression model that is specified as follows:
@@
Our choice of the graph Laplacian is designed to exploit information in the network to inform the model and force the predicted values to be smooth with respect to the topology of the line graph. That is, nodes are assigned values based on their connectivity in the network and adjacent nodes will have similar predicted values. In fact, the log prior can be expanded as
@@
showing that the term penalizes the weighted sum of squared differences for coefficients of adjacent vertices in the network (Kolaczyk and Csárdi, 2014). Moreover, pairs of nodes with larger edge weights, corresponding to more similar generalized lineups, will be forced to have more similar predicted values.
Given that $L_{w}$ is a symmetric matrix, it realizes the eigendecomposition $L_{w}=\Phi\Xi\Phi$, and we can therefore represent $\beta$ using a basis expansion with respect to the eigenvectors of $L_{w}$ by letting $\beta=\Phi\theta$ (Upton and Carvalho, 2024). In short, we use the eigenvectors of the Laplacian as predictor variables, find a set of corresponding coefficients $\hat{\theta}$ through Bayesian linear regression, and then perform matrix-vector multiplication to solve for the vertex-indexed intercepts $\widehat{\beta_{v}}=\widehat{Y_{v}}.$ We do not have to use the full set of eigenvectors to represent $\beta$. Instead, we can extract most of the information in $L_{w}$ through a low-order representation such that $\beta=\Phi_{1:\tau}\theta$, where $\Phi_{1:\tau}$ represents the first $\tau$ eigenvectors (ordered by the eigenvalues $\xi_{1}<\dots<\xi_{\tau}$). In practice, we select $\tau$ based on an ‘elbow’ method, examining sequential differences in the size of ${\xi_{i}}$.
To gain some intuition regarding LAPM, consider player C in the line graph in Figure 2. Player C is represent by its own node, which is also connected to nodes BC, AC, CD, and CE (representing pairs) and ABC and CDE (representing full lineups) because player C is part of those generalized lineups. Thus, the LAPM predicted value for C will be impacted, through the prior incorporating the Laplacian, by the predicted values/performance of all the aforementioned nodes.